Aftab tells his daughter, “Seven years ago, I was seven times as old as you were then. Also, three years from now, I shall be three times as old as you will be.”  Represent this situation algebraically and graphically.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Here I will provide you the algebraic representation :

Let the present age of Aftab be ‘x’.
And, the present age of his daughter be ‘y’.
Now, we can write,
seven years ago, Age of Aftab = x - 7
Age of his daughter = y - 7
According to the question,
x−7 = 7(y−7)
⇒ x−7 = 7y−49
⇒ x−7y = −42 ………………………(i)
Also, three years from now or after three years, Age of Aftab will become = x + 3.
Age of his daughter will become = y + 3
According to the situation given,
x+3 = 3(y+3)
⇒x+3 = 3y+9
⇒x−3y = 6 …………..…………………(ii)
This is required equation .
we can solve it further to get the values of x and y.

Thanks